OpenStudy (anonymous):

Solve for a: . 4a/9 =100/45 A. 5 B. 25 C. -5 D. -25

OpenStudy (anonymous):

5

OpenStudy (callisto):

4a/9 =100/45 a = (100/45) / (4/9) = 5